OpenStudy (zyberg):

What are integer solutions to this equation: x^4 - y^4 = 671?(Verify my solution, please)

OpenStudy (hollowdensity):

@Zyberg is that all?

OpenStudy (zyberg):

x^4 - y^4 = (x^2 - y^2)(x^2 + y^2); 671 = 11 * 61 (prime factors) since x^2 - y^2 < x^2 + y^2 x^2 - y^2 = 11 (x - y)(x + y)= 11 x - y = 1, since 11 is prime and, if x + y = 1, x or y would be <= 0. From that, x = y + 1; 2y + 1 = 11; y = 5, x = 6 Since a^2b works same as modulus, we get that answer pairs (6, 5) and (-6, -5) Is it right or are there more answers to this?

OpenStudy (zyberg):

@HollowDensity yes, it is everything that the problem states.

OpenStudy (kainui):

Yeah, looks rock solid to me.